OpenStudy (anonymous):

You deposit $2000 in Account A, which pays 2.25% annual interest compounded monthly. You deposit another $2000 in Account B, which pays 3% annual interest compounded monthly. When is the sum of the balance in both accounts at least $5000?

candycove (candycove):

Nice username.

OpenStudy (jdoe0001):

hmmm I assume you've used the compound interest formula already?

OpenStudy (anonymous):

yup. Im on the chapter using logarithms and exponential functions but it doesnt seem to work I know the answer is supposed to be 8.5 if that helps. I just dont know how to do it

OpenStudy (jdoe0001):

hmm

OpenStudy (anonymous):

So basically can you tell me the formulas in compounded interest form Note: make x in years, so in Account B you'll need to put x/12

OpenStudy (anonymous):

my formula was: 2000(1+0.0225/12)^12x + 2000(1+0.03/12)^12x is greater than or equal to 5000

OpenStudy (jdoe0001):

the cycle is yearly, so the cycle or "n = 1" in this case so \(\bf A=P\left(1+\frac{r}{1}\right)^{1t}\implies A=P(1+r)^t\) so.. the two balances need to be greater than or equal to 5000 so pretty much \(\bf 5000 \ge A + A\)

OpenStudy (jdoe0001):

lemme post what I have so far

OpenStudy (anonymous):

ok thanks i thought since it said compounded monthly it meant n=12 ... thats what we have been doing in class

OpenStudy (jdoe0001):

ohmmm shoot.. lemme reread that

OpenStudy (jdoe0001):

rats, yes, I read yearly smokes... yes ...so n = 12 ok

OpenStudy (anonymous):

\[2000(1+0.0225/12)^{12x} + 2000(1+0.03/12)^{12x}\] Ah yes I think I;ve got it so take the log of both sides, ignore the greater than make it just equals for now\[2000(1+0.0225/12)^{12x} + 2000(1+0.03/12)^{12x}=5000\]

OpenStudy (anonymous):

Take the log or ln or whatever you want of both sides which makes it\[\log_{10} (2000(1+0.0225/12)^{12x} + 2000(1+0.03/12)^{12x})\]

OpenStudy (anonymous):

Sorry \[\log10(2000(1+0.0225/12)^{12x}+2000(1+0.03/12)^{12x})=\log_{10}5000 \]

OpenStudy (anonymous):

Ahhh wait no can't do that

OpenStudy (anonymous):

lemme show you what my teacher did in class

OpenStudy (anonymous):

Wait let's restart\[2000(1+0.0225/12)^{12x} + 2000(1+0.03/12)^{12x}=5000\] Divide both sides by 2000 you get\[(12.0225)^{12x}+(1.03)^{12x}=2.5\]

OpenStudy (anonymous):

OHH Ok I see what your teacher did ALthough back up two steps that way takes you backward

OpenStudy (anonymous):

alrighty then :)

OpenStudy (anonymous):

Wait no Didn't see that x Hmm

OpenStudy (anonymous):

Ok yeah I honestly have no clue Either route I take mine or your teachers I'm at a dead end

OpenStudy (anonymous):

Sorry :/

OpenStudy (anonymous):

lol thats ok, even my teacher couldnt figure it out thanks for the effort though!!

OpenStudy (kropot72):

The problem reduces to the following, where t is the time in years: \[\large 1.001875^{12t}+1.0025^{12t}=2.5\] A value of t = 8.75 years is a close solution.

OpenStudy (anonymous):

yeah i got that reduced equation but how did you come to that answer? did u just do guess and check?

OpenStudy (kropot72):

I calculated the value of both terms on the left starting with 12t = 100. It soon emerges that12t = 105 gives a good approximation.

OpenStudy (anonymous):

alright thanks!

OpenStudy (kropot72):

You're welcome :)

OpenStudy (jdoe0001):

checking this or that way.... I don't see how to get the logs though to tweeze out the "t"

OpenStudy (anonymous):

yeah any way i try it the t is still in the exponent somewhere

OpenStudy (anonymous):

Yeah pretty sure the only ways to get an answer here is guessing and checking or graphing